OpenStudy (anonymous):

A toaster is plugged into an 120-volt outlet. If it has a resistance of 460 ohms, how much power does it use? A. 1,800 watts B. 3.8 watts C. 31 watts D. 55,000 watts E. 340 watts

OpenStudy (anonymous):

Power is generally given by the formula \[P = IV\] but by utilizing Ohm's law.. \[R = \frac{ V }{ I } \rightarrow I=\frac{ V }{ R }\] Plugging this into the original power equation we have \[P = (\frac{ V }{ R }) V = \frac{V^{2}}{R}\] Now just use the values you know for Voltage and Resistance!

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

You're welcome! Power can also be rearranged to just include current and resistance, so just be aware of that!

OpenStudy (anonymous):

Okay (:

OpenStudy (anonymous):

Do you get my other question ?